Light with a frequency of 6.85 × 1014 Hz lies in the violet region of the visible spectrum. What is the wavelength of this frequency of light?
1
Expert's answer
2020-02-21T07:53:19-0500

"\nu = 6.85 × 10^{14} Hz"

"\lambda =c/ \nu" , where

"c=3*10^8m\/s" --- speed of light

"\implies \lambda = (3*10^8)/(6.85*10^{15})"

"\implies \lambda = 4.38*10^{-8}=43.8nm"